# Metal-Poor Stars in the MW Disk: Resonant Cooling of Vertical Oscillations of Halo Stars in Barred Galaxies

### v1

@@ 4 Discussion
We have shown that a population of halo particles that oscillate about the potential midplane, i.e., about the Laplace surface, can decrease their amplitudes of oscillations in a short time period, $\sim 2-3$ oscillations. In other words, they are ‘cooled down’ abruptly. Other DM halo particles cool down but over longer time periods. When associating the DM particle trajectories with those of the halo metal-poor stars, one deals with the migration of halo stars to the disk. A number of interesting corollaries emerge as a result of this new process. But first of all, what is the underlying mechanism of this migration?
We start with an analogy: imagine a population of ping-pong balls above a moving floor. Both the balls and the floor can oscillate only along the vertical coordinate $z$, each with a different frequency and amplitude. We assume that the ball-floor collisions are elastic and that the ball population is released at various heights above $z=0$. Now, we analyze the results of the ball-floor collisions.
Table 1: Statistics of cooled halo particles for P00 and P60 models at $t=10$ Gyr and of cooled halos stars at $z=0$. $f_{\mathrm{cool}}$(abrupt) is the probability to find an abruptly cooled halo ‘star’ in the stellar disk moving with the disk kinematics; $N_{\mathrm{total}}$ is the number of halo particles with $\langle z_{\mathrm{max}}\rangle>2z_{\mathrm{disk}}$ from the entire sample, $N_{\rm total}$, at $t=0.5$ Gyr; prograde/retrograde ratio of cooled down halo particles found in the disk. $N_{*,\mathrm{cool}}$(total) is the estimated number of halo stars cooled down to the disk plane, using the Milky Way stellar halo.
@@
Stellar halo $N_{*,\mathrm{cool}}$(total) $9\times 10^{6}$ $5\times 10^{6}$
Some balls will acquire or lose kinetic energy from the collision with the floor depending on their velocity differences with the floor at the time of contact. Increasing or decreasing vertical kinetic energy is analogous to ‘heating up’ or cooling the balls, respectively. Some collisions will result in a substantial cooling of the vertical motions — this sub-population of balls will be trapped close to the floor. Some of the balls will be trapped for a prolonged time period, some will be re-heated and re-cooled multiple times. These outcomes demonstrate the possibility of the resonance cooling of the balls.
This toy model of ping pong balls – moving floor interactions can serve as a simplified analogy to vertical oscillations of halo stars and the horizontal motion of the stellar bar in the embedded galactic disk and their energy exchange. Halo stars can exchange energy and angular momentum with the spinning bar depending on their relative position before and after crossing the bar potential. As a result, the amplitude of the vertical oscillations will increase or decrease. A resonance between halo orbits and the bar perturbations leads to a fast energy transfer, i.e., during 2–3 vertical oscillations and cooling of these motions. This is the essence of the mechanism operating in barred galaxies and leading to the formation of the ‘cold’ population of metal-poor stars in the disk.
@@
where $N_{\rm cool}$ given by Figure 4. Results for both models are given in Table 1. $f_{\mathrm{cool}}\sim 4.7\%$ in P00, and $f_{\mathrm{cool}}\sim 2.9\%$ in P60.
The above results summarize the statistics of DM halo particles that experience the ‘cooling’ effect over various cooling intervals. To single out the extreme cooling effect, which proceeds over only 0.5 Gyr, we display the probability of such an abrupt event in Table 1. An example of such an abrupt cooling has been demonstrated in Figure 3a.
Observing the evolution of $N_{\rm cool}(t)$ in the top frame of Figure 4, a trend emerges that it is correlated with the amplitude evolution of the stellar bar. In both P00 and P60 models, this number increases until the buckling, i.e., $t\sim 3$ Gyr in P00, and $t\sim 4$ Gyr in P60. After the buckling, the bar regrows in the P00 model until the end of the simulation, as is $N_{\rm cool}(t)$. However, the remaining oval distortion in the P60 model after buckling (Figure 1) leads to a flattened $N_{\rm cool}(t)$ in P60.
@@
Hence, the total stellar mass within the equipotential surface $\Phi_{\mathrm{max}}$ in our model, and outside the disk with $|z|>6$ kpc, corrected for including only the bound stars, can be estimated as $M_{\rm*,total}\sim 1.3\times 10^{8}\,\mathrm{M_{\odot}}$. Consequently, using the same cooling probability $f_{\rm cool}$ in eq. (1) for the halo stars as for the DM, we obtain $N_{\rm*,cool}\sim 6\times 10^{6}$ in the P00 model, and $4\times 10^{6}$ in the P60 model (Table 1). To summarize, we expect the above $N_{\rm*,total}$ residing in the contemporary MW disk.
Figure 7: Distribution of the specific angular momentum, $j_{\rm z}$, from cooled particles of $N_{\mathrm{cool}}(t=10\,\mathrm{Gyr})$ (see Figure 4) for the P00 and P60 models. The vertical dashed lines show the average values for each distribution.
@@
To obtain the ratio of prograde-to-retrograde orbits of the cooled stellar halo population, we assume that the halo stars kinematics follows that of the DM, which can be wrong for the P60 model when applied to the MW — the inner stellar MW halo discussed here has been found to be essentially nonrotating (e.g., Freeman, 2012; Das et al., 2016; Deason et al., 2017). But because the properties of the inner stellar halos can be diverse, when applied to other galaxies, we still include this model. Under these conditions, we have calculated the averaged angular momentum over the last orbit before the end of the run (Figure 7). The prograde orbits are those spinning in the direction of the disk rotation. The resulting prograde-to-retrograde ratio of these particles is $1.2:1$ in the P00 model and is $0.8:1$ in the P60 model, when summed over all the cooled particles. However, these average ratios have little significance because the ratio depends on the radius, as we discuss below.
First, we find that the prograde and retrograde particles are distributed differently in the $xy$-plane at the end of the run (Fig. 8). In the P00 model, most of the prograde particles are distributed in a dumbbell shape with a major axis perpendicular to the major axis of the stellar bar. This dumbbell distribution reaches the radius beyond the bar size radius. In contrast, most retrograde particles dwell within the bar size radius and are distributed in an axisymmetric fashion. In the P60 model, with only an oval distortion at the end, the prograde particles have not shown any preferred azimuthal orientation. They are distributed more axisymmetrically than the prograde particles in the P00 model and extend to the region outside the bar size radius as well. The retrograde particles are distributed similarly to those in the P00 model. Most of them are within the bar size radius and distributed in an axisymmetric fashion.
Figure 8: The distribution of cooled DM particles of $N_{\mathrm{cool}}$($t=10$ Gyr) (see Figure 4) in the $x$-$y$ plane separated into prograde (left column) and retrograde (middle column) particles, for the P00 (top) and the P60 (bottom) models. The right column is part of randomly sampled disk particles showing the orientation of the stellar bar at the same time. The surface density of particles is plotted using contours with a linearly scaled color palette. The prograde and retrograde orbits have been identified by their instantaneous angular momentum $J_{\rm z}$.
@@
Second, we checked the radial dependence of prograde-to-retrograde ratio (Fig. 9). Outside the bar-size radius, in the P00 model, this ratio increases monotonically and becomes larger than 10 outside the corotation radius. In the P60 model, this ratio is not monotonic but reaches $\sim 2-2.5$ outside the corotation, then drops to unity, following a substantial increase at larger radii. Therefore, faster spinning stellar halos produce a more shallow distribution of prograde-to-retrograde orbits outside the corotation.
Unfortunately, no data exists for the radial distribution of the prograde-to-retrograde ratio of metal-poor stellar population in the disk. Recent observations, which indicate this ratio being about 3:1 for Gaia DR3 (Bellazzini et al., 2023) and other samples (e.g., Reimers, 1990; Beers et al., 1999; Starkenburg et al., 2017; Sestito et al., 2020; Lucey et al., 2021), may agree with both of our models. But detailed observations of the MW disk can provide the radial distribution of metal poor stars, and observations of other disk galaxies can reveal the scatter in this distribution which probably depends on the assembling history of their stellar halos.

### v2

@@ 4 Discussion
We have shown that population of halo particles that oscillate about the Laplace surface, can decrease their amplitudes of oscillations in a short time period, $\sim 2-3$ oscillations. In other words, they are ‘cooled down’ abruptly. Other DM halo particles cool down but over longer time periods. When associating the DM particle trajectories with those of the halo metal-poor stars, one deals with the migration of halo stars to the disk. A number of interesting corollaries emerge as a result of this new process. But first of all, what is the underlying mechanism of this migration?
We start with an analogy: a population of ping-pong balls above a moving floor. Both the balls and the floor can oscillate only along the vertical coordinate $z$, each with a different frequency and amplitude. We assume that the ball-floor collisions are elastic and that the ball population is released at various heights above $z=0$. Now, we analyze the results of the ball-floor collisions.
Table 1: Statistics of cooled halo particles for P00 and P60 models at $t=10$ Gyr and of cooled halo stars at redshift $z=0$. $N_{\mathrm{total}}$ is the number of sampled DM particles with $\langle z_{\mathrm{max}}\rangle>2z_{\mathrm{disk}}$ from the entire sample at $t=0.5$ Gyr; $f_{\mathrm{cool}}$(abrupt) is the probability of finding an abruptly cooled halo ‘star’ in the disk, following the disk kinematics; $f_{\mathrm{cool}}$(total) is the probability of finding a cooled halo ‘star’ in the disk due to all types of coolings; P/R is the prograde-to-retrograde ratio of cooled DM particles found in the disk. $N_{*,\mathrm{cool}}$(total) is the estimated number of halo stars cooled to the disk (see text), using the MW stellar halo.
@@
Stellar halo $N_{*,\mathrm{cool}}$(total) $1.2\times 10^{7}$ $7.6\times 10^{6}$
Balls will acquire/lose kinetic energy from the collision with the floor, depending on their velocity differences with the floor at the bouncing time. Increasing/decreasing vertical kinetic energy is analogous to ‘heating up’ or ‘cooling’ the balls. Some collisions will result in a substantial cooling of the vertical motions — this sub-population of balls will be trapped close to the floor. Some of the balls will be trapped for a prolonged time period, some will be re-heated and re-cooled multiple times. These outcomes demonstrate the possibility of the resonance cooling of the balls.
This toy model of ping pong balls – moving floor interactions can serve as a simplified analogy to vertical oscillations of halo stars and the horizontal motion of the stellar bar in the embedded galactic disk and their energy exchange. Halo stars can exchange energy and angular momentum with the spinning bar depending on their relative position before and after crossing the bar. As a result, the amplitude of the vertical oscillations will increase or decrease. A resonance between halo orbits and the bar perturbations leads to a fast energy transfer, i.e., during 2–3 vertical oscillations and cooling of these motions, especially because in 3-D potential the chaotic web related to resonances connects different phase space regions. This is the essence of the mechanism operating in barred galaxies and leading to the formation of the ‘cold’ population of metal-poor stars in the disk.
@@
where $N_{\mathrm{cool}}$ given by Figure 4. Results for both models are given in Table 1. $f_{\mathrm{cool}}\sim 4.7\%$ in P00, and $f_{\mathrm{cool}}\sim 2.9\%$ in P60.
The above results summarize the statistics of DM halo particles that experience the ‘cooling’ effect over various cooling intervals. To single out the extreme cooling effect, which proceeds over only 0.5 Gyr, we display the probability of such an abrupt event, $f_{\mathrm{cool}}$(abrupt), in Table 1. An example of such an abrupt cooling has been demonstrated in Figure 3a.
Observing the evolution of $N_{\mathrm{cool}}(t)$ in the top frame of Figure 4, a trend emerges that it is correlated with the amplitude evolution of the stellar bar. In both P00 and P60 models, this number increases until the buckling, i.e., $t\sim 3$ Gyr in P00, and $t\sim 4$ Gyr in P60. After the buckling, the bar regrows in the P00 model until the end of the simulation, as is $N_{\mathrm{cool}}(t)$. However, the remaining oval distortion in the P60 model after buckling (Figure 1) leads to a flattened $N_{\mathrm{cool}}(t)$ in P60.
@@
So is the cooling mechanism a resonant one? We easily find that for some DM particles the abrupt cooling correlates with the azimuthal angle between the particle and the stellar bar — when the particle crosses the Laplace plane slightly ahead of the bar having the same azimuth. This points to a resonant effect, when the energy of the vertical oscillation is transferred to the stellar bar. More generally, we rely on the structure of the phase space, as discussed in the section 1. The abrupt cooling is impossible if one thinks in terms of the regular phase space, ignoring the properties of the chaotic web in 3-D non-integrable potentials. The chaotic web does present many options of these transitions. Our search for cooling particles is a selective one, because we are interested in this specific aspect of evolution. Disk particles heating up into the halo could also be investigated, as well as other transitions between the distinct dynamical particle groups.
@@
Hence, the total stellar mass within the equipotential surface $\Phi_{\mathrm{max}}$ in our model, and outside the disk with $|z|>6$ kpc, including only the bound stars, can be estimated as $M_{\mathrm{*,total}}\sim 2.6\times 10^{8}\,\mathrm{M_{\odot}}$ — this was calculated exactly as $N_{\mathrm{total}}$ for the sampled DM particles, explained in section 3. Consequently, using the same cooling probability $f_{\mathrm{cool}}$ in eq. (1) for the halo stars as for the DM, we obtain $N_{\mathrm{*,cool}}\sim 1.2\times 10^{7}$ in the P00 model, and $7.6\times 10^{6}$ in the P60 model (Table 1). Hence, we expect the above $N_{\mathrm{*,cool}}$ residing in the contemporary MW disk.
Figure 7: Distribution of the specific angular momentum, $j_{\mathrm{z}}$, from cooled particles of $N_{\mathrm{cool}}(t=10\,\mathrm{Gyr})$ (see Figure 4) for the P00 and P60 models. The vertical dashed lines show the average values for each distribution.
@@
To obtain the ratio of prograde-to-retrograde orbits of the cooled stellar halo population, we assume that the halo star kinematics follows that of the DM, which is questionable for the P60 when applied to the MW — the inner stellar MW halo discussed here has been found to be essentially nonrotating (e.g., Freeman, 2012; Das et al., 2016; Deason et al., 2017). Under these conditions, we calculate the instantaneous angular momentum, $J_{z}$, of cooled particles. The prograde orbits are those spinning with the disk. The resulting prograde-to-retrograde ratio, P/R, of these particles is $1.2:1$ in the P00 model and $0.8:1$ in the P60 model. However, these average ratios have little significance because the ratio depends on the radius, as we discuss below.
First, we find that the prograde and retrograde particles are distributed differently in the $xy$-plane at the end of the run (Fig. 8). In the P00 model, most of the prograde particles are distributed in a dumbbell shape with a major axis perpendicular to the major axis of the stellar bar. This dumbbell distribution reaches the radius beyond the bar-size radius. In contrast, most retrograde particles dwell within the bar-size radius and are distributed in an axisymmetric fashion. In the P60 model, with only an oval distortion at the end, the prograde particles have not shown any preferred azimuthal orientation. They are distributed more axisymmetrically than the prograde particles in the P00 model and extend to the region outside the bar-size radius as well. The retrograde particles are distributed similarly to those in the P00 model. Most of them are within the bar-size radius and distributed in an axisymmetric fashion.
Figure 8: The distribution of cooled DM particles of $N_{\mathrm{cool}}$($t=10$ Gyr) (see Figure 4) in the $x$-$y$ plane separated into prograde (left column) and retrograde (middle column) particles, for the P00 (top) and the P60 (bottom) models. The right column is part of randomly sampled disk particles showing the orientation of the stellar bar at the same time. The surface density of particles is plotted using contours with a linearly scaled color palette. The prograde and retrograde orbits have been identified by their instantaneous angular momentum $J_{\mathrm{z}}$.
@@
Second, we checked the radial dependence of the P/R. Outside the bar radius, in the P00 model, this ratio increases monotonically and becomes larger than 10 outside the corotation radius. In the P60 model, this ratio is not monotonic but reaches $\sim 2-2.5$ outside the corotation, then drops to unity, following a substantial increase at larger radii. Therefore, faster spinning stellar halos produce a more shallow distribution of P/R orbits outside the corotation.
No data exists for the radial distribution of the P/R for metal-poor stellar population in the disk. Observations indicating that P/R being $\sim 3:1$ for Gaia DR3 (Bellazzini et al., 2024) and other samples (e.g., Reimers, 1990; Beers et al., 1999; Starkenburg et al., 2017; Sestito et al., 2020; Lucey et al., 2021) may agree with our models. Detailed observations of the MW can provide the radial distribution of metal-poor stars, and observations of other disk galaxies can reveal the scatter in this distribution which probably depends on the assembling history of their stellar halos.
In summary, we have presented a novel mechanism of cooling the vertical oscillations of halo stars in barred galaxies with nonspinning and spinning DM halos. Using numerical simulations, we argue that this mechanism is associated with a chaotic transition between regular regions separated by the chaotic phase space generated by the strong bar perturbation. And the chaotic web does allow a large choice of possible transitions between the regular islands of phase space. Finally we remark that the particle cooling presented here is not contradicting the conservative nature of the $N$-body systems, because globally the kinetic energy lost by the particle subsample is transferred to the rest of the system. This energy transfer, more generally the action, would be much slower if the system were close to integrable, like spherical systems.